Find the slope, if possible, of the line passing through each pair of points.
(2,-1) and (6,3)

Found 2 solutions by Maths68, lenny460:
Answer by Maths68(1474)

Slope of the straight line = m =(y1-y2)/(x1-x2)
(x1,y1)=(2,-1)
(x2,y2)=(6,3)
m =(y1-y2)/(x1-x2)
m=(-1-3)/(2-6)
m=(-4)/(-4)
m=1
Slope of the straight line = m = 1

Answer by lenny460(1073)

After plotting the straight line graph of:
(2,-1)
(6,3)
The Slope = rise/run = 4/4 = 1
The other method:
(y2 - y1)/(x2 - x1)
3 - (-1)/6 -2
3 - (-1) = 3 + 1 = 4
6 -2 = 4
4/4 = 1

or
-1 - 3/2 - 6
-1 - 3 = -4
2 - 6 = -4
-4/-4 = 4/4 = 1
Therefore:
The Slope - m = 1


obuong3@aol.com